# Uncovering Cultural Representation Disparities in Vision-Language Models

### v1

@@ 6 Conclusion
The results we obtained contradict the popular assumption of western bias in VLMs, The Open-Ended experiments yielded results that bias toward few popular countries and non necessarily western nations as seen in Figure 18. Further at the task of distinguishing and recognizing cultures among other similar countries through a MCQ task, the models performed nearly as good as the rest of the world in most cases barring a few sparsely populated countries and a few from Central-Africa as seen in Figure 19. These findings were consistent across multiple runs of the same setup, across languages as well as across all of the models tested.
6.1 Likely Cause for contrasting findings
The key difference between the samples in our datasets and those from the other works is the complexity of images provided. In most of the images, we used it is difficult to understand and predict the country of origin without a strong vision ability. For example, the Afghanistan example from Figure 31 is difficult to classify without spotting the hint of Afghan flag from one of the people from the image. While the appearance and ethnicity of the person on the left may mislead the VLM. This challenging nature of a large portion of images used could have lead to the contrasting findings in some instances.
6.2 Scope for further analysis
Unlike analyses on LLMs, or those works over categorizing language resource levels at a language level, it is difficult to analyze to what extent the pre-training and instruction-tuning dataset distribution might have impacted the biases found. Although some previous works list country wise distribution of some open datasets (Kuznetsova et al., 2020) (Shankar et al., 2017), it is not known how the newer models’ dataset distributions might have varied. This would require more open-data models.
@@ Appendix C Reproducibility
Inference was done through Cohere’s API for Aya-Vision-8B and Aya-Vision-32B through the default hyperparameters with a seed value of 1024. The rest of the models were used through OpenRouter’s API through the default hyper-parameters with a seed value of 1024. The experiments were repeated thrice and the overall accuracy varied between 1-1.2%, with some countries’ accuracy varying up to 1.5%.

### v2

@@ 7 Conclusion
We investigated cultural biases in Vision-Language Models using a diverse country-level image classification task. Our findings show that biases are not uniformly Western but instead reflect overrepresentation of certain countries in training data. Model performance varied across prompt types, languages, image features, and perturbations—highlighting limitations in robustness and cultural generalization.
These results call for greater transparency in dataset composition and the need for more culturally inclusive evaluation methods to ensure fairer and more globally representative VLMs.
@@ Appendix C Reproducibility
Inference was done through Cohere’s API for Aya-Vision-8B and Aya-Vision-32B through the default hyperparameters with a seed value of 1024. The rest of the models were used through OpenRouter’s API through the default hyper-parameters with a seed value of 1024. The experiments were repeated thrice and the overall accuracy varied between 1-1.2%, with some countries’ accuracy varying up to 1.5%. The costs associated with all experiments combined were 850\$ through OpenRouter and 250\$ Cohere API credits. The experiments were run on TPUs costing 0.35\$/hr with the costs reaching  60\$.
Category Description
Appearance (Attire) Attires of some people from the image, clothes being hanged in the background, etc.
Appearance (People) Appearance / visual perception of people’s ethnicity, presence of any celebrities, etc.
Architecture (Exterior) Building facades, monuments, bridges, outdoor structures, and any external architectural elements visible in the scene.
Architecture (Interior) Indoor environments e.g. rooms, corridors, staircases, furniture, and interior design details.
Landscape (Water) Bodies of water such as oceans, rivers, lakes, waterfalls, ponds, and any aquatic scenery.
Landscape (Air) Aerial / bird’s-eye views, landscapes captured from above, clouds, sky scenes, and horizon vistas.
Landscape (Vegetation) Forests, grasslands, gardens, crops, shrubs, foliage patterns, plant life, or visible greenery.
Texts/Scripts/Posters Signs, banners, billboards, labels, handwritten or printed text, posters, and any other written or graphic messaging.
Patterns/Designs Decorative motifs, surface textures, fabric prints, wallpaper or tile patterns, abstract designs, and repetitive graphical elements.
Table 3: Overview of the image categories used to analyse model performance as a function of the type of image.
@@ Appendix G Country wise accuracies in each experimental setting
The accuracies obtained over samples of each country through each experimental setup can be seen in Table 4 to Table 8.
@@
Figure 35: A sample from our dataset and its corresponding response (GPT-4o-Mini)